OpenStudy (anonymous):

Which of the following best defines 2 to the power of 2 over 3? Cube root of 4 Square root of 4 Cube root of 2 Square root of 2

geerky42 (geerky42):

2 to the power of 2 over 3 \(\Large = 2^{2/3} = \sqrt[3]{2^2}\)

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

@geerky42 i dont understand

geerky42 (geerky42):

Do you understand that part: 2 to the power of 2 over 3 \(\Large =2^{2/3}\) ?

OpenStudy (anonymous):

yes

geerky42 (geerky42):

well, \(\Large 2^{2/3} = \left(2^2\right)^{1/3} = \sqrt[3]{2^2}\) if you have fraction exponent, you can rewrite it as root, for example: \(\Large x^{1/n} = \sqrt[n]{x}\) does that make sense?

OpenStudy (anonymous):

yea

geerky42 (geerky42):

so which is answer?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

d i mean

geerky42 (geerky42):

? do you know what \(\large 2^2\) is?

OpenStudy (anonymous):

4

geerky42 (geerky42):

yes so \(\Large \sqrt[3]{2^2}=\sqrt[3]{4}\) Do you know what \(\Large \sqrt[3]{~~}\) is called?

OpenStudy (anonymous):

cubed

geerky42 (geerky42):

you are thinking of \(\Large ~^3\). \(\Large\sqrt[3]{~~}\) is called cube root.

geerky42 (geerky42):

So \(\Large\sqrt[3]{4}\) can be read as "Cube root of 4" is that clear?

OpenStudy (anonymous):

ok yea very clear

OpenStudy (anonymous):

thx